OpenStudy (anonymous):

Find the slope and the y-intercept of the equation x − 4(y + 1) − 5 = 0.

OpenStudy (owlcoffee):

Try bringing it to the following form: \[y=mx+c\]

OpenStudy (anonymous):

How would i go about doing that.

OpenStudy (owlcoffee):

First do a distributive with that (-4) you have there, we'd end up with: \[x-4y-4-5=0\] simplifying: \[x-4y-9=0\] Now, all you have to do is leave the "y" alone in the left side of the equal sign.

OpenStudy (anonymous):

slope = −1/4 and y-intercept = −9/4 slope = 1/4 and y-intercept = −9/4 slope = 1/4 and y-intercept = 9/4 slope = −1/4 and y-intercept = 9/4 the answers,i dont know how to get from that to those

OpenStudy (anonymous):

-4y=-x+9 y=1/4x-9/4

OpenStudy (anonymous):

whats the slope whats the y intercept

OpenStudy (anonymous):

it didnt tell me that

OpenStudy (anonymous):

well it's simple there the formulais y=mx+b m being the slope and b being y-intercept